OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[\left( 4^{2} \right)^{6}\]

OpenStudy (anonymous):

4^12

OpenStudy (chriss):

\[(4^{6*2})=4^{12}\]

OpenStudy (anonymous):

only multiplic potences ;)

OpenStudy (turingtest):

\[\huge (4^2)^6=(2^2)^{2^6}=(2^2)^{2\cdot6}=2^{2\cdot2\cdot6}\]\[\huge =2^{24}\]

OpenStudy (turingtest):

if you want toe get it completely down to scratch...